### v1

@@ 6.1 Test problems and numerical setup
$\frac{\partial u}{\partial t}+\frac{\partial f(u)}{\partial x}$ $=0,\quad f(u)=\frac{u^{2}}{2},$ (B)
$\frac{\partial u}{\partial t}+\frac{\partial f(u)}{\partial x}$ $=0\quad f(u)=\frac{4u^{2}}{4u^{2}+(1-u)^{2}},$ (BL)
with periodic boundary conditions in space. Equation (B) is the Burgers equation and serves as the simplest nonlinear hyperbolic PDE, and (BL) is the Buckley–Leverett equation and can have much richer solution structure than Burgers equation owing to its non-convex flux $f$—see [45, Section 16.1.1] for detailed discussion on the solution structure of (BL). For both PDEs, we use the square wave initial condition $u_{0}(x)=1$ for $x\in(-0.5,0)$ and $u_{0}(x)=0$ for $x\in(-1,1)\setminus(-0.5,0)$. We choose this initial data because it gives rise to both rarefaction and shock waves. We solve (B) on the time domain $t\in[0,4]$, and (BL) on $t\in[0,2]$; solution plots are shown in Fig. 2. Note that for (BL), when the LLF flux (6) with (10) is used, reconstructions are always limited to be within the physical range $u\in[0,1]$: Any $u_{i+1/2}^{\pm}>1$ are mapped to $u_{i+1/2}^{\pm}=1$, and any $u_{i+1/2}^{\pm}<0$ are mapped to $u_{i+1/2}^{\pm}=0$; see Remark SM2.1 in the Supplementary Materials for further discussion.
@@
Figure 2: Solution plots for numerical test problems: Burgers equation (B) (left column) and Buckley–Leverett equation (BL) (right column). Plots in the top row are space-time contours, with the black lines being 20 contour levels evenly spaced between 0 and 1. Plots in the bottom row are solution cross-sections at the times indicated in the legends. For the Burgers equation, the initial discontinuity at $x=-0.5$ propagates as a rarefaction wave with speed $1$ while the one at $x=0$ propagates as a shock wave with speed $1/2$. At $t=1$ these two waves merge and propagate at a reduced speed thereafter (see Supplementary Material Section SM7 for further details). For the Buckley–Leverett equation, both discontinuities propagate as compound waves: They contain both a shock and an attached rarefaction.
@@
All tests use nonlinear F-relaxation with CF-splitting factor $m=8$ (Line 2 in Algorithm 1). In all tests, a single MGRIT iteration is used to approximately solve each linearized problem (Line 7 in Algorithm 1), with the initial MGRIT iterate set equal to the corresponding right-hand side vector of the linearized system, i.e., the current nonlinear residual. For simplicity, we consider only two-level MGRIT, and perform a single F-relaxation for the MGRIT pre-relaxation using $m=8$. This MGRIT initialization strategy is equivalent, although computationally cheaper, to using a zero initial MGRIT iterate followed by pre CF-relaxation on the first MGRIT iteration on the fine level. [^5] For $\bm{r}_{k}\approx\bm{0}$, the linearization error is small, and, thus, the linearized problem is a highly accurate model for the nonlinear residual equation (27). Thus, for sufficiently small $\bm{r}_{k}$, our relaxation and initial iterate strategy is effectively the same as if no nonlinear relaxation were applied and MGRIT was applied to the linearized problem with FCF-relaxation. Note that FCF-relaxation is typically used in the context of MGRIT for linear hyperbolic problems, with F-relaxation sometimes not sufficient for obtaining a convergent solver [16].
The approximate truncation error correction at coarse-grid time steps, i.e., the matrix inversion in (58), is applied exactly via LU factorization. In our previous work for linear problems [16, 14], and for the linear results in Section SM5, these linear systems were approximately solved using a small number of GMRES iterations. Numerical tests (not shown here) indicate that GMRES is possibly less effective for the current linearized problems, and we hope to address this in future work.
@@ 6.2 Results
If Algorithm 1 is to be effective when approximate solves of the linearized systems are used, it must converge adequately in the best-case setting that direct solves are used for the linearized systems (that is, when Line 5 is executed rather than Line 7). Numerical tests confirming the efficacy of Algorithm 1 in the special case of direct linear solves can be found in Supplementary Material Section SM2.
@@
Solving to residual tolerances as tight as those in Figs. 3 and 4 is important for demonstrating stability of the algorithm [^6] Typically, when applied to advection-dominated problems, multigrid-in-time methods are divergent on a set of error modes when used with naive direct coarse-grid discretizations. If these error modes are not present or only weakly present in the initial iterate, the method will appear to converge on initial iterations, but will diverge on later iterations due to exponential growth of these problematic error modes [17, 16, 28]. and understanding the mesh-dependence of the asymptotic convergence rate. However, such tight tolerances likely do not result in more accurate approximations to the underlying PDE solution. For shocked scalar problems, the discretizations we consider in this work do not converge in the $L^{\infty}$-norm, and their convergence rate in the discrete $L^{1}$-norm is, at best, ${\cal O}(h)$ for strictly convex fluxes, and more generally is ${\cal O}(\sqrt{h})$ [33, 62, 38]. Thus, the practical value of iterating until the algebraic residual is point-wise small globally is questionable. Section SM7 considers a Burgers problem with $n_{x}=1024$, showing that the approximation after just two nonlinear iterations is as good an approximation to the true PDE solution as the sequential time-stepping solution is.
@@ 6.3 Speed-up potential
We again stress that our chief goal in this paper is developing and verifying a new parallel-in-time strategy for difficult nonlinear problems, and that this alone represents significant work. With this is mind, we now briefly discuss the potential of our method to provide speed-up over sequential time-stepping when it is implemented in parallel. In our previous work for linear problems [16, 14], when solving to tight residual tolerances, we achieved speed-up factors on the order of two to 12 times, depending on the problem. We anticipate our method developed here for nonlinear problems is capable of generating similar or slightly smaller speed-ups provided a fully multilevel MGRIT method can be used (i.e., not just two levels), and that the truncation error correction can be approximated with a cheap iterative method (i.e., not LU factorization). Our reasoning is as follows. The nonlinear convergence rates here are comparable to the linear convergence rates in [16, 14]. We note also that a comparable amount of work per iteration is done on the fine grid in each approach. Here we apply nonlinear F-relaxation, a nonlinear residual evaluation at C-points, and an F-relaxation on the linearized problem, while in [16, 14] FCF-relaxation was used.
@@ 7 Conclusion and future outlook
This work leaves open many avenues for future research. An immediate next step is to develop a parallel implementation. In [15] we describe extending the methodology in this paper to the viscous Burgers equation.
@@
These Supplementary Materials are structured as follows. Section SM1 provides an introduction to the topic of polynomial reconstruction. Section SM2 presents further numerical tests. Section SM3 develops error estimates for the FV semi-Lagrangian method described in Section 5.3. Section SM4 develops error estimates for method-of-lines discretizations applied to a simplified linear conservation law. Section SM5 presents an MGRIT solver for the linear conservation law (cons-lin) when it is discretized (on the fine grid) with standard linear method-of-lines discretizations. Section SM6 develops an expression for the ideal truncation error in the coarse-grid operator (58) where the fine-grid problem is a linearized method-of-lined discretization. Section SM7 considers the issue of over-solving the discretized problem for a Burgers problem.
@@
In this section, we give a brief overview of polynomial reconstruction. Polynomial reconstruction is a key ingredient of the method-of-lines discretization described in Section 2, and of the semi-Lagrangian discretization discussed in Section 5.3. Moreover, the error estimates we develop for these discretizations (see Sections SM4 and SM3) require some understanding of this reconstruction process. More detailed descriptions on polynomial reconstruction can be found in the excellent works by Shu [57, 58].
@@
$S^{\ell}=\{{\cal I}_{-\ell},\ldots,{\cal I}_{r}\},$ (SM1)
in which the left-most cell is ${\cal I}_{-\ell}$, and the right-most cell is ${\cal I}_{r}$. In this notation, $\ell,r\geq 0$, and we have that $\ell+r+1=k$. We say that $S^{\ell}$ has a left shift of $\ell$. The number of cells $k$ determines the order of accuracy of the reconstruction. For example, a 1st-order reconstruction uses the stencil $S^{0}=\{{\cal I}_{0}\}$, i.e., $\ell=0$, $r=0$, and a (centered) 3rd-order reconstruction uses the stencil $S^{1}=\{{\cal I}_{-1},{\cal I}_{0},{\cal I}_{1}\}$, i.e., $\ell=r=1$. Note the cells of $S^{\ell}$ in (SM1) cover the interval $x\in[x_{-\ell-1/2},x_{r+1/2}]$.
Now, we seek a polynomial $q_{k-1}(x)\approx u(x)$, $x\in[x_{-\ell-1/2},x_{r+1/2}]$, of degree at most $k-1$ such that it recovers the cell averages of the function $u(x)$ over the $k$ cells in the stencil $S$:
$\frac{1}{h}\int_{x_{j-1/2}}^{x_{j+1/2}}q_{k-1}(x)\operatorname{d}\!x=\frac{1}{h}\int_{x_{j-1/2}}^{x_{j+1/2}}u(x)\operatorname{d}\!x=:\bar{u}_{j},\quad\textrm{for}\quad j=-\ell,\ldots,r.$ (SM2)
@@
$U(x)=\int_{-\infty}^{x}u(\zeta)\operatorname{d}\!\zeta,\quad\textrm{and}\quad U^{\prime}(x)=u(x).$ (SM3)
@@
$\bar{u}_{j}:=\frac{1}{h}\int_{x_{j-1/2}}^{x_{j+1/2}}u(x)\operatorname{d}\!x=\frac{1}{h}\left(U(x_{j+1/2})-U(x_{j-1/2})\right),\quad\textrm{for}\quad j=-\ell,\ldots,r.$ (SM4)
@@
$Q_{k}(x_{j+1/2})=U(x_{j+1/2}),\quad\textrm{for}\quad j=-\ell-1,\ldots,r.$ (SM5)
@@
$\bar{u}_{j}$ $:=\frac{1}{h}\left(U(x_{j+1/2})-U(x_{j-1/2})\right)=\frac{1}{h}\left(Q_{k}(x_{j+1/2})-Q_{k}(x_{j-1/2})\right),$ (SM6)
$=\frac{1}{h}\int_{x_{j-1/2}}^{x_{j+1/2}}Q_{k}^{\prime}(x)\operatorname{d}\!x,$ (SM7)
@@
$U(x)-Q_{k}(x)=\frac{1}{(k+1)!}\prod\limits_{j=-\ell-1}^{r}\big{(}x-x_{j+1/2}\big{)}U^{(k+1)}(\varphi(x)),\quad x\in[x_{-\ell-1/2},x_{r+1/2}],$ (SM9)
@@
Overall, we find that the nonlinear solver converges quickly, including with mesh-independent iteration counts in many cases, even for non-differentiable discretizations. Note that we have also tested the solver on Burgers problems in which a shock forms from smooth initial conditions (not shown here), and the solver then appears to, in general, converge faster than for the Burgers problem considered in Fig. 2.
@@
SM3 Error estimate for semi-Lagrangian method and approximate truncation error

### v2

@@ 6.1 Test problems and numerical setup
(B) $\frac{\partial u}{\partial t}+\frac{\partial f(u)}{\partial x}$ $=0,\quad f(u)=\frac{u^{2}}{2},$
(BL) $\frac{\partial u}{\partial t}+\frac{\partial f(u)}{\partial x}$ $=0,\quad f(u)=\frac{4u^{2}}{4u^{2}+(1-u)^{2}},$
with periodic boundary conditions in space. Equation (B) is the Burgers equation and serves as the simplest nonlinear hyperbolic PDE, and (BL) is the Buckley–Leverett equation and can have much richer solution structure than Burgers equation owing to its non-convex flux ff—see [45, Section 16.1.1] for detailed discussion on the solution structure of (BL). For both PDEs, we use the square wave initial condition $u_{0}(x)=1$ for $x\in(-0.5,0)$ and $u_{0}(x)=0$ for $x\in(-1,1)\setminus(-0.5,0)$. We choose this initial data because it gives rise to both rarefaction and shock waves. Additional numerical tests using a smooth initial condition can be found in Supplementary Materials Section SM3, where solver convergence is qualitatively similar to that for the square wave initial condition. We solve (B) on the time domain $t\in[0,4]$, and (BL) on $t\in[0,2]$; solution plots are shown in Fig. 2. Note that for (BL), when the LLF flux (6) with (8) is used, reconstructions are always limited to be within the physical range $u\in[0,1]$: Any $u_{i+1/2}^{\pm}>1$ are mapped to $u_{i+1/2}^{\pm}=1$, and any $u_{i+1/2}^{\pm}<0$ are mapped to $u_{i+1/2}^{\pm}=0$; see Supplementary Materials Remark SM2.1 for further discussion.
@@
Figure 2: Solution plots for numerical test problems: Burgers equation (B) (left column) and Buckley–Leverett equation (BL) (right column). Plots in the top row are space-time contours, with the black lines being 20 contour levels evenly spaced between 0 and 1. Plots in the bottom row are solution cross-sections at the times indicated in the legends. For the Burgers equation, the initial discontinuity at x=0.5x=-0.5 propagates as a rarefaction wave with speed 11 while the one at $x=0$ propagates as a shock wave with speed $1/2$. At $t=1$ these two waves merge and propagate at a reduced speed thereafter (see Supplementary Material Section SM9 for further details). For the Buckley–Leverett equation, both discontinuities propagate as compound waves: They contain both a shock and an attached rarefaction.
@@
All tests use nonlinear F-relaxation with CF-splitting factor $m=8$ (Line 2 in Algorithm 1). In all tests, a single MGRIT iteration is used to approximately solve each linearized problem (Line 7 in Algorithm 1), with the initial MGRIT iterate set equal to the corresponding right-hand side vector of the linearized system, i.e., the current nonlinear residual. We perform a single F-relaxation for the MGRIT pre-relaxation using a coarsening factor of $m=8$. This MGRIT initialization strategy is equivalent, although computationally cheaper, to using a zero initial MGRIT iterate followed by pre CF-relaxation on the first MGRIT iteration on the fine level. [^5] For $\bm{r}_{k}\approx\bm{0}$, the linearization error is small, such that the linearized problem is a highly accurate model for the nonlinear residual equation (13). Thus, for sufficiently small $\bm{r}_{k}$, our relaxation and initial iterate strategy is effectively the same as if no nonlinear relaxation were applied and MGRIT was applied to the linearized problem with FCF-relaxation. Note that FCF-relaxation is typically used in the context of MGRIT for linear hyperbolic problems, with F-relaxation sometimes not sufficient for obtaining a convergent solver [15]. For simplicity, we consider only two-level MGRIT; however, the interested reader can find multilevel results for first-order discretizations in Supplementary Materials Section SM4, where the multilevel algorithm proceeds analogously to those in our previous works using modified semi-Lagrangian coarse operators [41, 14, 15].
The approximate truncation error correction at coarse-grid time steps, i.e., the matrix inversion in (40), is applied exactly via LU factorization. In our previous work for linear problems [15, 14], and for the linear results in Supplementary Materials Section SM7, these linear systems were approximately solved using a small number of GMRES iterations. Numerical tests (not shown here) indicate that GMRES is possibly less effective for the current linearized problems, and we hope to address this in future work.
@@ 6.2 Results
If Algorithm 1 is to be effective when approximate solves of the linearized systems are used, it must converge adequately in the best-case setting that direct solves are used for the linearized systems (that is, when Line 5 is executed rather than Line 7). Numerical tests confirming the efficacy of Algorithm 1 in the special case of direct linear solves can be found in Supplementary Materials Section SM2.
@@
Solving to residual tolerances as tight as those in Figs. 3 and 4 is important for demonstrating stability of the algorithm [^6] Typically, when applied to advection-dominated problems, multigrid-in-time methods are divergent on a set of error modes when used with naive direct coarse-grid discretizations. If these error modes are not present or only weakly present in the initial iterate, the method will appear to converge on initial iterations, but will diverge on later iterations due to exponential growth of these problematic error modes [16, 15, 27]. and understanding the mesh-dependence of the asymptotic convergence rate. However, such tight tolerances likely do not result in more accurate approximations to the underlying PDE solution. For shocked scalar problems, the discretizations we consider in this work do not converge in the $L^{\infty}$-norm, and their convergence rate in the discrete $L^{1}$-norm is, at best, ${\cal O}(h)$ for strictly convex fluxes, and more generally is ${\cal O}(\sqrt{h})$ [32, 62, 37]. Thus, the practical value of iterating until the algebraic residual is point-wise small globally is questionable. Supplementary Materials Section SM9 considers a Burgers problem with $n_{x}=1024$, showing that the approximation after just two nonlinear iterations is as good an approximation to the true PDE solution as the sequential time-stepping solution is.
@@ 6.3 Speed-up potential
We again stress that our chief goal in this paper is developing and verifying a new parallel-in-time strategy for difficult nonlinear problems, and that this alone represents significant work. With this is mind, we now briefly discuss the potential of our method to provide speed-up over sequential time-stepping when it is implemented in parallel. In our previous work for linear problems [15, 14], when solving to tight residual tolerances, we achieved speed-up factors on the order of two to 12 times, depending on the problem. We anticipate our method developed here for nonlinear problems is capable of generating similar or slightly smaller speed-ups provided that the truncation error correction can be approximated with a cheap iterative method (i.e., not LU factorization). Our reasoning is as follows. The nonlinear convergence rates here are comparable to the linear convergence rates in [15, 14]. We note also that a comparable amount of work per iteration is done on the fine grid in each approach. Here we apply nonlinear F-relaxation, a nonlinear residual evaluation at C-points, and an F-relaxation on the linearized problem, while in [15, 14] FCF-relaxation was used.
@@ 7 Conclusion and future outlook
This work leaves open many avenues for future research. An immediate next step is to develop a parallel implementation; extensions to problems in multiple spatial dimensions are also important, and are likely to be based on the multi-dimensional MGRIT methodology developed in [14]. In [42] we discuss the extension of this work to one-dimensional hyperbolic systems, including the acoustic equations and the Euler equations of gas dynamics.
Acknowledgments
Critical feedback from anonymous referees is gratefully acknowledged.
@@ Acknowledgments
These Supplementary Materials are structured as follows. Section SM1 provides an introduction to the topic of polynomial reconstruction. Sections SM2, SM3, and SM4 present further numerical tests. Section SM5 develops error estimates for the FV semi-Lagrangian method described in Section 5.3. Section SM6 develops error estimates for method-of-lines discretizations applied to a simplified linear conservation law. Section SM7 presents an MGRIT solver for the linear conservation law (cons-lin) when it is discretized (on the fine grid) with standard linear method-of-lines discretizations. Section SM8 develops an expression for the ideal truncation error in the coarse-grid operator (40) where the fine-grid problem is a linearized method-of-lined discretization. Section SM9 considers the issue of over-solving the discretized problem for a Burgers problem.
@@
In this section, we give a brief overview of polynomial reconstruction. Polynomial reconstruction is a key ingredient of the method-of-lines discretization described in Section 2, and of the semi-Lagrangian discretization discussed in Section 5.3. Moreover, the error estimates we develop for these discretizations (see Sections SM6 and SM5) require some understanding of this reconstruction process. More detailed descriptions on polynomial reconstruction can be found in the excellent works by Shu [57, 58].
@@
(SM1) $S^{\ell}=\{{\cal I}_{-\ell},\ldots,{\cal I}_{r}\},$
[…]
Now, we seek a polynomial $q_{k-1}(x)\approx u(x)$, $x\in[x_{-\ell-1/2},x_{r+1/2}]$, of degree at most k1k-1 such that it recovers the cell averages of the function $u(x)$ over the $k$ cells in the stencil $S$:
(SM2) $\frac{1}{h}\int_{x_{j-1/2}}^{x_{j+1/2}}q_{k-1}(x)\operatorname{d}\!x=\frac{1}{h}\int_{x_{j-1/2}}^{x_{j+1/2}}u(x)\operatorname{d}\!x=:\bar{u}_{j},\quad\textrm{for}\quad j=-\ell,\ldots,r.$
@@
(SM3) $U(x)=\int_{-\infty}^{x}u(\zeta)\operatorname{d}\!\zeta,\quad\textrm{and}\quad U^{\prime}(x)=u(x).$
@@
(SM4) $\bar{u}_{j}:=\frac{1}{h}\int_{x_{j-1/2}}^{x_{j+1/2}}u(x)\operatorname{d}\!x=\frac{1}{h}\left(U(x_{j+1/2})-U(x_{j-1/2})\right),\quad\textrm{for}\quad j=-\ell,\ldots,r.$
@@
(SM5) $Q_{k}(x_{j+1/2})=U(x_{j+1/2}),\quad\textrm{for}\quad j=-\ell-1,\ldots,r.$
@@
(SM6) $\bar{u}_{j}$ $:=\frac{1}{h}\left(U(x_{j+1/2})-U(x_{j-1/2})\right)=\frac{1}{h}\left(Q_{k}(x_{j+1/2})-Q_{k}(x_{j-1/2})\right),$
(SM7) $=\frac{1}{h}\int_{x_{j-1/2}}^{x_{j+1/2}}Q_{k}^{\prime}(x)\operatorname{d}\!x,$
@@
(SM9) $U(x)-Q_{k}(x)=\frac{1}{(k+1)!}\prod\limits_{j=-\ell-1}^{r}\big(x-x_{j+1/2}\big)U^{(k+1)}(\varphi(x)),\quad x\in[x_{-\ell-1/2},x_{r+1/2}],$
@@
Refer to caption    Refer to caption
Refer to caption    Refer to caption
@@
Refer to caption    Refer to caption
Refer to caption    Refer to caption
@@
Overall, we find that the nonlinear solver converges quickly, including with mesh-independent iteration counts in many cases, even for non-differentiable discretizations. Note that we have also tested the solver on Burgers problems in which a shock forms from smooth initial conditions (not shown here), and the solver then appears to, in general, converge faster than for the Burgers problem considered in Figure 2.
SM3 Additional numerical tests: Different initial condition
@@
Figure SM4: Additional numerical tests for Burgers equation (B) using a smooth initial condition. Top row: Space-time contours, with the black lines being 20 contour levels evenly spaced between 0 and 1 (left); cross-sections of the solution at the times indicated in the legend (right). Bottom row: Residual convergence histories for the nonlinear solver Algorithm 1 applied to a 1st-order discretization (left), and a 3rd-order discretization (right). Solid lines correspond to linear systems being approximately solved by one two-level MGRIT iteration, and broken lines correspond to linear systems being solved exactly.
In this section we consider additional numerical tests to complement those from Section 6.1, considering a smooth initial condition in contrast to the square wave that was used there. Specifically, we use the initial condition $u_{0}(x)=0.5+0.4\cos(\pi x)+0.1\sin(3\pi x)$}. The numerical setup for these tests is the same as described in Section 6.1. We again consider the Burgers equation (B) and the Buckley–Leverett equation (BL). We solve (B) on the time domain $t\in[0,4]$, and on a sequence of space-time meshes with resolutions $(n_{x},n_{t})\in\{(64,78),(128,309),(256,617),(512,1233),(1024,2465),(2048,4929),\\ (4096,9857)\}$; solution plots are shown in the top row of Figure SM4. We solve (BL) on the time domain $t\in[0,2]$, and on a sequence of space-time meshes with resolutions $(n_{x},n_{t})=\{(64,188),(128,375),(256,748),(512,1494),(1024,2986),(2048,5971),\\ (4096,11941)\}$; solution plots are shown in the top row of Figure SM5.
For each problem we consider both 1st- and 3rd-order discretizations, but consider only the LLF flux, and for the 3rd-order WENO discretization we consider only the Newton linearization (c.f., Figures 3 and 4 where the GLF flux was considered and also Picard linearization of the WENO discretization).
Residual convergence histories for (B) are shown in the bottom row of Figure SM4, and those for (BL) are shown in the bottom row of Figure SM5. Considering the convergence histories, we see largely the same trends as occurred for the square wave initial condition considered in Section 6.1; see Figures 3 and 4. That is, when direct solves are used for the linearized space-time problems, the solver converges in a small number of iterations with mesh-independent convergence rates. When the linearized space-time problems are solved inexactly using a single two-level MGRIT iteration, there seems to be some slight deterioration in convergence speed in most cases; however, in all cases, the nonlinear iteration counts appear to be bounded as the space-time mesh is refined. Moreover, iteration counts for this initial condition are qualitatively similar to those for the square wave considered in the main paper.
@@
Figure SM5: Additional numerical tests for Buckley–Leverett equation (BL) using a smooth initial condition. Top row: Space-time contours, with the black lines being 20 contour levels evenly spaced between 0 and 1 (left); cross-sections of the solution at the times indicated in the legend (right). Bottom row: Residual convergence histories for the nonlinear solver Algorithm 1 applied to a 1st-order discretization (left), and a 3rd-order discretization (right). Solid lines correspond to linear systems being approximately solved by one two-level MGRIT iteration, and broken lines correspond to linear systems being solved exactly.
SM4 Additional numerical tests: Multilevel MGRIT
Since there are many complicated details in the main paper, when employing MGRIT to approximately solve the linearized systems we used only a two-level method for simplicity’s sake. Here we discuss the extension of this MGRIT method to the multilevel setting, relying on recursively applying the two-level method, analogous to our multilevel extensions of modified semi-Lagrangian schemes in prior works [14, 15, 41].
To keep things as simple as possible, we consider multilevel extensions for first-order discretizations only, i.e., $p=k=1$. The methodology can be extended to the third-order discretizations considered in the main paper, but it is more complicated because the truncation-error correction uses weighted stencils, as in the $k=2$ case shown in (41).
Some of the notation from the main paper is now slightly changed since we need to account explicitly for the level index $\ell\in\mathbb{N}_{0}$, with $\ell=0$ denoting the finest level. Let us write $\Phi^{(n)}_{\ell}$ as the MGRIT time-stepping operator that steps from $t_{n}\to t_{n}+m^{\ell}\delta t$, since on level $\ell$ the time-step size is $m^{\ell}\delta t$. Let ${\cal S}^{(n,m^{\ell}\delta t)}$ be a first-order semi-Lagrangian discretization taking one step of size $m^{\ell}\delta t$ to step from $t_{n}\to t_{n}+m^{\ell}\delta t$. Moreover, let $\bm{\varepsilon}_{\ell}^{n}\in\mathbb{R}^{n_{x}}$ be the vector of mesh-normalized departure points that the semi-Lagrangian method ${\cal S}^{(n,m^{\ell}\delta t)}$ uses to advance the solution from $t_{n}$ to $t_{n}+m^{\ell}\delta t$.
The level $\ell=1$ time-stepping operator from the main paper is given by (40), and for first-order discretizations it simplifies as
(SM10) $\Phi_{1}^{(n)}$ $=\Big(I+{\cal T}_{\mathrm{ideal},1}^{n}-{\cal T}_{\mathrm{direct},1}^{n}\Big)^{-1}{\cal S}^{(n,m\delta t)}$
(SM11) $=\Big(I+{\cal D}_{1}\operatorname{diag}(\bm{\delta}^{n}_{1}){\cal D}_{1}^{\top}\Big)^{-1}{\cal S}^{(n,m\delta t)},\quad\bm{\delta}^{n}_{1}=h^{2}g_{2}(\bm{\varepsilon}^{n}_{1})+\sum_{j=0}^{m-1}\bm{\beta}^{n+j},$
in which the coefficients $\bm{\delta}^{n}_{1}$ are obtained from plugging in ${\cal T}_{\mathrm{ideal},1}^{n}$ from (41), and ${\cal T}_{\mathrm{direct},1}^{n}$ from (48). That is, $\bm{\delta}^{n}_{1}$ accounts approximately for the difference in the leading-order truncation error terms between the “ideal” and “direct” methods used to cover the interval $t_{n},t_{n}+m\delta t]$, with the former being $m$ linearized method-of-lines steps of size $\delta t$, and the latter being one semi-Lagrangian step of size $m\delta t$.
In the forthcoming derivation, on coarser levels $\ell>1$, we approximate $m$ smaller semi-Lagrangian steps with a single large semi-Lagrangian step using the truncation-error-based approximation (40). That is, on levels $\ell>1$ we invoke the approximation (where the product is assumed to expand leftwards with increasing $k$)
(SM12) $\prod_{k=0}^{m-1}{\cal S}^{(n+k,m^{\ell-1}\delta t)}\approx\Big(I+{\cal T}_{\mathrm{ideal},\ell}^{n}-{\cal T}_{\mathrm{direct},\ell}^{n}\Big)^{-1}{\cal S}^{(n,m^{\ell}\delta t)},$
where now the “ideal” method on the interval $t_{n},t_{n}+m^{\ell}\delta t]$ is $m$ semi-Lagrangian steps of size $m^{\ell-1}\delta t$, and the “direct” method is one semi-Lagrangian step of size $m^{\ell}\delta t$. The right hand side of (48) gives the general form of our approximation of the semi-Lagrangian truncation error. Plugging this in gives, for $\ell>1$,
(SM13) ${\cal T}_{\mathrm{ideal},\ell}^{n}-{\cal T}_{\mathrm{direct},\ell}^{n}$ $=\sum_{k=0}^{m-1}-h^{2}{\cal D}_{1}\operatorname{diag}\big(g_{2}(\bm{\varepsilon}_{\ell-1}^{n+k})\big){\cal D}_{1}^{\top}+h^{2}{\cal D}_{1}\operatorname{diag}\big(g_{2}(\bm{\varepsilon}_{\ell}^{n})\big){\cal D}_{1}^{\top},$
(SM14) $={\cal D}_{1}\operatorname{diag}(\bm{\Delta}_{\ell}^{n}){\cal D}_{1}^{\top},\quad\bm{\Delta}_{\ell}^{n}:=h^{2}\Big[g_{2}(\bm{\varepsilon}_{\ell}^{n})-\sum_{k=0}^{m-1}\big(g_{2}(\bm{\varepsilon}_{\ell-1}^{n+k})\big)\Big].$
That is, $\bm{\Delta}_{\ell}^{n}$ accounts approximately for the difference in the leading-order truncation error terms between using $m$ semi-Lagrangian steps of size $m^{\ell-1}\delta t$ to cover the interval $t_{n},t_{n}+m^{\ell}\delta t]$ vs. one semi-Lagrangian step of size $m^{\ell}\delta t$.
To generalize the level $\ell=1$ operator (SM10) to the multilevel setting, we now consider the ideal operator on level $\ell=2$ and the following sequence of approximations to it:
(SM15) $\Phi_{{{\mathrm{ideal},2}}}^{(n)}$ $:=\Phi^{(n+m-1)}_{1}\,\cdots\,\Phi^{(n+1)}_{1}\Phi^{(n)}_{1}$,
(SM16) $=\prod_{k=0}^{m-1}\Big[\Big(I+{\cal D}_{1}\operatorname{diag}(\bm{\delta}^{n+k}_{1}){\cal D}_{1}^{\top}\Big)^{-1}{\cal S}^{(n+k,m\delta t)}\Big],$
(SM17) $\approx\left[\prod_{k=0}^{m-1}\Big(I+{\cal D}_{1}\operatorname{diag}(\bm{\delta}^{n+k}_{1}){\cal D}_{1}^{\top}\Big)^{-1}\right]\left[\prod_{k=0}^{m-1}{\cal S}^{(n+k,m\delta t)}\right],$
(SM18) $\approx\left[\prod_{k=0}^{m-1}\Big(I+{\cal D}_{1}\operatorname{diag}(\bm{\delta}^{n+k}_{1}){\cal D}_{1}^{\top}\Big)^{-1}\right]\left[\Big(I+{\cal D}_{1}\operatorname{diag}(\bm{\Delta}^{n}_{2}){\cal D}_{1}^{\top}\Big)^{-1}{\cal S}^{(n,m^{2}\delta t)}\right],$
(SM19) $\approx\left[I+{\cal D}_{1}\operatorname{diag}\left(\bm{\Delta}^{n}_{2}+\sum_{k=0}^{m-1}\bm{\delta}^{n+k}_{1}\right){\cal D}_{1}^{\top}\right]^{-1}{\cal S}^{(n,m^{2}\delta t)}:=\Phi^{(n,m^{2}\delta t)}_{2}$.
We now explain the above approximations. The first is that the $m$ semi-Lagrangian operators commute with the $m$ truncation error terms. The next approximation is to replace $m$ semi-Lagrangian steps of size $m\delta t$ by one single semi-Lagrangian step of size $m^{2}\delta t$ and to include the leading-order truncation error terms. The final approximation has two steps: The first is to replace the $m$ products with a sum (analogous to the usual approximation one makes for MGRIT applied to a diffusion problem [20]), and the second is to pull the two matrices under the same inverse and neglect higher-order terms from their product; see [14, 15, 41].
Based on (SM10) and (SM19) we propose the following coarse-level operator on level $\ell>0$:
(SM20) $\Phi^{(n,m^{\ell}\delta t)}_{\ell}=\Big[I+{\cal D}_{1}\operatorname{diag}(\bm{\sigma}^{n}_{\ell}){\cal D}_{1}^{\top}\Big]^{-1}{\cal S}^{(n,m^{\ell}\delta t)},\quad\ell\in\mathbb{N},$
where the coefficient $\bm{\sigma}^{n}_{\ell}$ is defined recursively by
(SM21) $\bm{\sigma}^{n}_{\ell}=\begin{cases}\bm{\delta}^{n}_{1},\quad&\ell=1,\\[4.30554pt] \bm{\Delta}^{n}_{\ell}+\sum\limits_{k=0}^{m-1}\bm{\sigma}^{n+k}_{\ell-1},\quad&\ell>1.\end{cases}$
Now we present multilevel numerical results using the coarse-level operator (SM20), extending several of the two-level results from Section 6 of the main paper. To generate a multievel hierarchy, we recursively coarsen in time by a factor of $m=8$ until doing so would result in a temporal coarse mesh with fewer than two points. The finest temporal resolution for the Burgers problem is $n_{t}=10241$ and for the Buckley–Leverett problem it is $n_{t}=11941$; in either case, the MGRIT hierarchy for each problem has 5 levels. We use one MGRIT V-cycle per inner iteration, analogous to how one two-level iteration was used in the main paper. Our MGRIT relaxation strategy is level dependent, using F-relaxation on the finest MGRIT level, as for the two-level solves used in the main text, and then FCF-relaxation on all coarser levels.
Numerical results are shown in Fig. SM6, with multilevel results shown in the top row, and the bottom row showing the two-level results from the bottom row of Figure 3 from the main paper. For the Burgers problem (left) there is no deterioration in convergence when moving from two-level to multilevel. For the Buckley–Leverett problem there is a slight deterioration in iteration count for the finest resolution problems, where one or two more iterations are required to reach the halting tolerance.
Refer to caption    Refer to caption
Refer to caption    Refer to caption
Figure SM6: Residual convergence histories for the nonlinear solver Algorithm 1 applied to 1st-order accurate discretizations using the LLF flux. Left column: Burgers (B). Right column: Buckley–Leverett (BL). Top row: Linearized systems are solved approximately with one MGRIT V-cycle. Bottom row: Solid lines correspond to linear systems being approximately solved by one two-level MGRIT iteration, and broken lines correspond to linear systems being solved exactly.
SM5 Error estimate for semi-Lagrangian method and approximate truncation error